OpenStudy (fouzberzerk):

Estimate the value of this expression by rounding each number to the nearest whole unit. (1.94)(7.51)(3.74) A. 56 B. 64 C. 42 D. 21

OpenStudy (fouzberzerk):

@kimbllse777

OpenStudy (anonymous):

(2)(8)(4) so 2×8=16×4=64

OpenStudy (fouzberzerk):

Great, thanks for the quick responses.

OpenStudy (anonymous):

You're welcome :)

OpenStudy (tkhunny):

Show us YOUR work on the next one.